FUSE Mission Status Report

Mission Status Report #7      Star Date: June 30, 1999

Photo: Science Team Lead, Bill Oegerle accepts the "key to FUSE" from Jim Moore, NASA/GSFC Director of Flight Projects, at Launch Day ceremony at JHU. (Photo courtesy of NASA/GSFC photographer Debbie McCallum). (Click photo to see larger version.) See more JHU Launch Day Photos.


Date: Wed, 30 Jun 1999 15:35:00 -0400 (EDT) 
From: Bill Oegerle 
Subject: FUSE IOC, June 30

FUSE IOC news, June 30, 1999 

Today was a very full day for FUSE, and this will be a longer report
than normal.  A number of us (Dave Sahnow, Scott Friedman and I) were up
at the crack of dawn (actually 3:30 am EDT) to watch the turn-on of
Far-UV detector #2.  Low voltage was turned on, and the stim pulsers
were activated.  Everything responded as planned, and all currents and
voltages were nominal.  Both detector segments were receiving the
expected count rates.  Over the next couple of orbits, the detector
temperatures steadily rose to expected values.  

Next, Jeff Kruk arrived for the activation of FES-B, which was booted at
5:30am EDT.  Again, all telemetry looked nominal.  FES-A will be turned
on a few days from now.  

An orbit later, we turned on Far-UV detector #1.  Again, everything looks

The real fun started when we performed the Bright Object Sensor BOS) 
calibration and performed a planned slew to beta=160 degrees (20 degrees 
from the sun!).  This worked great, and the baffle doors cooperated by 
staying closed (as they should--whew!).  The BOS triggered at approximately 
the correct angle from the sun, but we don't know the trigger angle 
precisely yet because the telemetry is on the spacecraft recorders, and 
will be downlinked to the ground later.  But all signs indicate that the 
BOS works as designed.  Kudos to the University of Colorado. 

But we weren't finished!  Next we performed a BOS bright earth check, to
make sure the bright object sensors would not trigger when we looked at
bright earth.  Again the BOS passed the test.  This test was done
after performing a slew of almost 180 degrees, which the spacecraft handled
nicely.  We are starting to get real comfortable with slews carried
out by the attitude control system.  Kudos to Orbital Sciences.

Later this afternoon, we will turn on the Focal Plane Assembly (FPA)
electronics. Ken Brownsberger (Univ of Colorado) is here to monitor the

Finally, time permitting, we will upload a new thermal control table.
The spectrograph temperature will be increased over the next 3-4 days
to a temperature of about 30 degrees (current temp is less than 10 deg).

Well, I think that is enough for one day.  Tomorrow we will be doing an
engineering checkout of FES-B and taking the first images with the FES
(baffle doors still closed) to determine the dark current in the CCD. 
We will then begin mapping the SAA.

We conclude with a note of appreciation to the flight ops team for
putting in lots of long hours getting the spacecraft and instrument
turned on and checked out. Today's award for exceptional duty goes to Tom
Jennings, who has been here for well over 12 hours now.  Tom's broad
knowledge of the instrument has been very valuable during this
check-out period.  Thanks, Tom!

Reported by: Bill Oegerle, Science Ops Lead

Return to the List of Status Reports